A nuclear power reactor generates 4.0 x 109 W of power. In one year (365.25 days), what is the change in the mass of the nuclear fuel due to the energy being taken from the reactor?
Answer #1

nuclear power reactor generates 4.0 x 109 W

one year (365.25 days),

4.0 * 10^9 W = 4.0 * 10^9 J/s
4.0* 10^9 J/s * 60 s/min * 60 min/hr * 24 hr/day * 365.25 days/year =1.26*10^17 J

change in energy = change in mass * speed of light^2

1.26*10^17 J = change in mass * (3*10^8)^2
change in mass = 1.4 kg